Anchor Ltd paid $15,000 last quarter for a feasibility study regarding the demand for motor-boat replacement parts which would require the purchase of a new metal-shaping machine. Today, they wish to conduct an analysis of the proposed project. The machine costs $250,000 and will operate for five years and tax rules allow the machine to be depreciated to zero over a five-year life. The machine is expected to produce sales of $135,000 annually for the five years. Anchor has already agreed to sell the machine in five years’ time to an unrelated firm for $80,000. The project will result in a $35,000 increase in accounts receivable and require an increase in inventory levels by $20,000 to $95,000. Anchor has negotiated with its bank to borrow $180,000 to help pay for the project. Loan repayments are $48,000 each year for five years. If Anchor buys the machine they will be able to use some equipment that they currently own. This is part of the driving force in the decision making as it enables the company to save money in not buying additional new equipment. This equipment was bought for $120,000 six years ago and could be sold today for $63,000. This equipment has been written off for tax purposes and would be worthless in five years’ time. What are the relevant cash flows for each year of the new machine’s life? Assume the company tax rate is 30%.
Answer #1
Cost of feasibility study($15000) is a sunk cost and not relevant
Initial Cash Flow (Year0)
Machine cost ($250,000)
Less Borrowing $180,000
A Net initial cash flow for machine ($70,000)
B Cash flow due to increase in accounts receivable ($35,000)
C Cash flow due to increase in inventory ($20,000)
D Opportunity cost of using existing equipment ($63,000)
E=A+B+C+D Total Initial Cash Flow ($188,000)
Pmt Loan Repayment per year $48,000
Nper Number of years                         5
Pv Amount of loan $180,000
RATE Interest Rate 10.424845% (Using RATE function of excel with Nper=5, Pmt=48000,Pv=-180000)
Excel Command : RATE(5,48000,-180000)
A B C=10.4248%*A D=B-C E=A-D
Year Beginning Loan balance Loan Repayment Interest Principal Ending Loan Balance
1 $180,000 $48,000 $18,765 $29,235 $150,765
2 $150,765 $48,000 $15,717 $32,283 $118,482
3 $118,482 $48,000 $12,352 $35,648 $82,833
4 $82,833 $48,000 $8,635 $39,365 $43,468
5 $43,468 $48,000 $4,532 $43,468 $0
ANNUAL CASH FLOWS:
Annual Loan Repayment ($48,000)
Interest Tax Shield: a b=a*30%
Year Interest Payment Interest tax shield
1 $18,765 $5,629
2 $15,717 $4,715
3 $12,352 $3,705
4 $8,635 $2,591
5 $4,532 $1,359
Depreciation Tax shield
c Annual Depreciation $50,000 (250000/5)
d=c*30% Annual Depreciation Tax shield $15,000
e Annual Before tax sales Revenue $135,000
f=e*(1-0.3) Annual After tax sales Revenue $94,500
After tax salvage value of machine=80000*(1-0.3) $56,000
YEAR 0 1 2 3 4 5
I Initial Cash Flow ($188,000)
F Annual after tax Sales Revenue $94,500 $94,500 $94,500 $94,500 $94,500
G Loan repayment ($48,000) ($48,000) ($48,000) ($48,000) ($48,000)
H Interest tax shield $5,629 $4,715 $3,705 $2,591 $1,359
J Depreciation tax shield $15,000 $15,000 $15,000 $15,000 $15,000
K $56,000
L=I+F+G+H+J+K Net Cash Flow ($188,000) $67,129 $66,215 $65,205 $64,091 $118,859